Let G be a nonabelian group of order 10 having golden ratio 1+√

5

2

as the neutral

element. Then:

(a) Verify class equation for G.

(b) Find all elements of Inn(G).

(c) Verify that G/Z(G) ∼= Inn(G).

(d) For some elements x, y, z ∈ G: verify the commutator identity: [xy, z] = [x, z][x, z, y][y, z]


1
Expert's answer
2021-07-23T09:13:16-0400

Part a.

Since the order of G is 10, so the golden ratio divides the order of G.

"=\frac{|G|}{1+\frac{\sqrt{5}}{2}}\\
=\frac{|10|}{1+\frac{\sqrt{5}}{2}}\\
=\frac{2*10}{1+\sqrt{5}}\\"

So, the class equation for G is

"[1+ \frac{\sqrt{5}}{2}]^n" where n = 1,...,10


Part b

In this part of the question, we will employ the general class equation for G.This will help us get the  Inn(G) elements (First four elements).

The  Inn(G) elements are given below

"[1+ \frac{\sqrt{5}}{2}]^n\\
[1+ \frac{\sqrt{5}}{2}]^0= [1]\\
[1+ \frac{\sqrt{5}}{2}]^1= [1+ \frac{\sqrt{5}}{2}]\\
[1+ \frac{\sqrt{5}}{2}]^2=[1+ \frac{\sqrt{5}}{2}]^2\\
[1+ \frac{\sqrt{5}}{2}]^3=[1+ \frac{\sqrt{5}}{2}]^3\\"

"[1],[1+ \n2\n5\n\u200b\n \n\u200b\n ],[1+ \n2\n5\n\u200b\n \n\u200b\n ] ^\n2\n ,[1+ \n2\n5\n\u200b\n \n\u200b\n ] ^\n3"


Part c

"=\frac{G}{Z(G)}\\
=\frac{10}{2}\\
=2"

and

"Inn(G)=\frac{10}{2}=5\\
So,\frac{G}{Z(G)}= Inn(G)."


Part d

From the question we are asked to verify the commutator identity.

"LHS\\
[xy,z]=xyz-zxy\\
RHS\\
[x, z][x, z, y][y, z] = x[y,z]+[x,z]y\\
\space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space = x[yz-zy]+[xz-zx]y\\
\space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space = xyz-xzy+xzy-zxy\\
\space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space = xyz-zxy\\"

The LHS and the RHS are equal. Hence the commutator is identical.